OpenStudy (tester97):

How do i find x? 87x+54(451+458)=70 @HamSandwich

OpenStudy (hamsandwich):

K, so, first, add the terms that are in the parentheses. You'll get 909.

OpenStudy (anonymous):

87x+54(451+458)=70 I will try to solve your equation. 87x+49086=70(Simplify both sides of the equation) 87x+49086+−49086=70+−49086(Add -49086 to both sides) 87x=−49016 (87x) 87 = (−49016) 87 (Divide both sides by 87) x=−563.402299

OpenStudy (hamsandwich):

way to copy and paste from wolframalpha, Joanna

OpenStudy (tester97):

You worked that way too fast joanna >.>

OpenStudy (hamsandwich):

then, after you get 909, multiply said 909 by 54 to get 49086. @JoannaPink, you spent literally 1 second typing that. I'm not that dumb.

OpenStudy (anonymous):

Online. I do online school, so I do the work, then copy and paste it

OpenStudy (anonymous):

I am fast with the keyboard

OpenStudy (hamsandwich):

So then, it's 87x + 49086 = 70. After that, subtract 49086 from both sides to get 87x = -49016. Joanna, you were also here for about 5 seconds before you posted your answer. Again, I'm not that dumb.

OpenStudy (hamsandwich):

The last step to find out x is dividing each side by 87. So, x = -563.40229885057471264367816091954. By the way, that doesn't happen. Stop lying.

OpenStudy (hamsandwich):

My 2005 emachines running windows xp with severe lag doesn't even take long.

OpenStudy (hamsandwich):

Plus, that would only take long for you. Not other people.

OpenStudy (hamsandwich):

then get out. Again, you're really doing a terrible job at ignoring.

OpenStudy (tester97):

Newb do you even block button?